Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $40 per unit. Variable expenses are $20.00Req 1 Req 2 Req3 Req 4A Req 4B Req 5A Req 5B Req 6 What is the products CM ratio? CM ratio %Use the CM ratio to determine the break-even point in dollar sales. (Do not round intermediate calculations.) Break-even poinIf this years sales increase by $45,000 and fixed expenses do not change, how much will net operating income increase? Net oAssume the president expects this years sales to increase by 10%. Using the degree of operating leverage from last year, whaWhat is the degree of operating leverage based on last years sales? (Round intermediate calculations and final answer to 2 dThe sales manager is convinced that a 11% reduction in the selling price, combined with a $64,000 increase in advertising, woThe sales manager is convinced that a 11% reduction in the selling price, combined with a $64,000 increase in advertising, woThe president does not want to change the selling price. Instead, he wants to increase the sales commission by $2.10 per unit
Answer #1
1) CM ratio
contribution margin/sales
540,000/1,080,000
0.5
50%
2) Break even sales in dollars
fixed cost/contribution margin ratio
180,000/50%
360000
3) Net operating income will increase by
45,000*50%
22500
4-a) Degree of operating leverage
contribution margin/net income
540,000/360,000
1.5
4b) percentage increase in net income
10*1.5
15 %
5)-a) unit sale = 1,080,000/40
27000
now unit sales =27000*125%
33750
selling price new = 40*89%=$35.6
sales        (33750*35.6) 1201500
Variable expense (33750*20) 675000
Contribution margin 526500
fixed expense (180,000+64000) 244000
Net operating income 282500
b) Decrease in net income 77500
per unit
6) sales 40
Variable expense 22.1
Contribution margin 17.9
increase in sales 25%
(33750*17.9) - (180,000+x) = 360,000
604125 - 180,000-x =340,000
x                                 = 84,125 asnwer